DIFFERENTIAL GEOMETRY/PDE SEMINAR

Monday, June 11 Padelford C-401

2:30 PM

An informal talk on 2-dimensional point vortices

Tadashi Tokeida

(Université du Québec à Montréal)

I will discuss how point vortices move, their exact solutions, where in nature we see them, and their relevance to some theorems in symmetric Hamiltonian systems.